OpenStudy (anonymous):

in which quadrant the system fall: y=x-1 y=-3x-5

OpenStudy (anonymous):

\[y = x- 1\] \[y = -3x -5\] so make them equal to each other \[x - 1 = -3x -5\] combine like-terms \[4x = -4\] dived by 4 \[x = -1\] So now plug into any equation so lets pick the first one... \[y = -1 - 1\] \[y = -2\] so now you have the point \[(-1,-2)\] and what quadrant is that in? Quadrent III

OpenStudy (jhannybean):

Good job @rnick !